40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 3 16:43:44 2011, Katrin, Update, Green Locking, SHG temperature (YARM) 
    Reply  Thu Nov 3 17:04:23 2011, Koji, Update, Green Locking, SHG temperature (YARM) 
       Reply  Thu Nov 3 17:24:59 2011, Zach, Update, Green Locking, SHG temperature (YARM) 
          Reply  Fri Nov 4 13:53:33 2011, Katrin, Update, Green Locking, SHG temperature (YARM) 
    Reply  Wed Nov 9 16:49:17 2011, kiwamu, Update, Green Locking, Y end laser temperature with slow input connected 
Message ID: 5797     Entry time: Thu Nov 3 16:43:44 2011     Reply to this: 5798   5852
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: SHG temperature (YARM) 

Plugging in the thermal feedback BNC cable to the laser reduced the DC voltage of the green PDH photo diode from 3.12 V to 1.5V off resonance.

The power emitted by the laser was the same as in the case without that cable. Note LT, i.e. measured crystal temperature, of the laser show a

different value when the BNC is connected, but the manual clearly states that this display does not work properly if a cable is connected to the

slow BNC plug, an offset is added.

The power of the 532nm light behind the SHG oven has been reduced from 1mW to 0.4mW. I changed the crystal temperature such that the power

of the green light is 1mW. With this new temperature setting the laser can be locked again.

 

  w/o BNC cable at slow plug w/ BNC cable at slow plug
T (°C) 29.776 29.776
LT (°C) 39.2 48.4
1064nm power (mW) 440 448
Temp. at TC200 (°C) 35.7 36.4
532nm power in front of Isolator (mW) 1.0 1.0

 

 

 

ELOG V3.1.3-